Mrs. Vessell's students created some beautiful tessellations and she wants to put them up in her room. If she has 12 tessellations and wants to line them up across the back of her room but only has room to put up 6, in how many different ways can she display the works of art?


1
Expert's answer
2021-06-06T15:43:26-0400

"N=C^6_{12}=\frac{12!}{6!6!}=924" ways